# Fast, Modular, and Differentiable Framework for Machine Learning-Enhanced Molecular Simulations

### v1

@@ III.3 Machine Learning/Molecular Mechanics
An alternative to describing the system fully using an MLIP is to employ a hybrid approach, where part of the system is described using an MLIP, and the remaining part by a classical force field. This is often called ML/MM, standing for machine learning/molecular mechanics, a term derived from QM/MM used to describe the combination of quantum mechanical ab initio calculations and molecular modelling. While the field of QM/MM is quite advanced, in particular in choosing a proper embedding approach for the electrostatic interactions [40], it is significantly less explored for the ML/MM approach. The current approaches are based on a mechanistic embedding, and as such the electrostatic interactions are treated on the MM level [41] or ignored. To define an ML/MM system, one defines a ml_system which is modeled with an MLIP and a mm_sytem described by a classical force field. In addition, one needs define ml_atoms, a list containing the atom indices modeled by the ml_system, which then results in the new combined system. Example code on how to achieve this is presented in Fig. 2(a).
III.4 Handling of Units
In DIMOS, we do not choose a specific unit system to perform all calculations, but rather directly set the necessary constants in the chosen unit system. This has the big advantage over established approaches, that the computations are very interpretable, which can be beneficial when testing against established simulation packages which may use different unit systems. In particular, for “unitless” simulations as often performed in statistical physics, this allows for an effortless comparison to many specialized code bases not using units. In most cases the units in DIMOS are adopted from the AMBER simulation suite, unless otherwise specified. Globally changing the unit system is straightforward and can be achieved by calling
dimos.init_constants_in_unit_system(unit_system)
where unit_system is the target unit system encoded as string, accepting both names of specific simulation packages or energy units.
III.5 Molecular Dynamics and Energy Minimization
@@
With the system description in place, we now present the implemented methods that update the position of the atoms, with various goals in mind.
III.5.1 Molecular Dynamics
@@
The basis of MD is numerically solving Newtons equation of motion given by $\frac{\partial E}{\partial\vec{r}}=-\vec{F}=-m\vec{a}$. To do so, we implement the ubiquitous methods by Verlet [29] in its velocity variant [42], which numerically integrates Newtons equation in time and hence conserve the total energy of the shadow Hamiltonian.
@@
Here, $\vec{x}_{i}$ is the position of atom $i$, $\vec{v}_{i}$ and $\vec{a}_{i}$ are the velocity and acceleration. Each step is understood to be performed for all atoms $i$. Since the updates are performed synchronously, the update for different atoms can easily be performed in parallel. In step 3. and 6., we have indicated that in case of constraints in the system, i.e., when certain distances or angles between atoms are kept constant, this should be performed at that place.
III.5.2 Constant Temperature Simulations
The most direct way to control the temperature of a system, i.e., to sample the canonical ensemble, are thermostats, applied in step 7.. For an overview on different thermostatting strategies, see Ref. [43] and references therein. We implement the Berendsen, Andersen, and Lowe-Andersen thermostat [44]. These thermostats work by either rescaling the atoms velocities or by picking new velocities from the Maxwell-Boltzmann distribution for a fraction of atoms or pairs of atoms [44].
It is also possible to define equations that deviate from Newton’s equation of motion, leading for example to the Langevin and Brownian dynamics equation [43]. These integrators require adapted integration schemes [45].
Another way to control the temperature to introduce additional degrees of freedom that act as coupling to an external heatbath. This modifies the integration, where the additional degree of freedoms are explicitly integrated. We implement the Nosé–Hoover and Nosé–Hoover chain dynamics [43], which introduce one or more additional harmonic ‘particles’ that couple the system’s temperature to the target temperature.
Practically, in DIMOS, one defines all different type of thermostats or ways to perform the dynamics as separate integrators. Below, the code necessary to define the Langevin and Nosé–Hoover chain dynamics, as well as dynamics using the velocity Verlet integrator in conjunction with the Andersen thermostat is demonstrated.
@@ III.5.2 Constant Temperature Simulations
III.5.3 Constant Temperature and Pressure Simulations
Finally, to reproduce many experimental conditions correctly, it is necessary to simulate the system at constant pressure. We implement an isotropic and anisotropic Monte Carlo barostat [46], which works by proposing new box sizes and accepting the new state with the usual Metropolis Monte Carlo criterion. In particular, one proposes new box sizes for the simulation cell, and then repositions the coordinates of the atoms within this box to correspond to the changed dimensions. A high acceptance rate is achieved by displacing the center of each molecule instead of treating each atom individually, as this way one avoids distortions within each molecule. Alternatively, in cases where molecules are not as clearly defined such as for MLIPs, it is also possible to simply displace every atom independently.
Following OpenMM [26], we have also implemented an option that adapts the proposal distribution heuristically so that the target acceptance rate lays within the user specified bounds. This avoids situations where either only small changes to the volume is proposed, which are then accepted very often, but lead only to small actual changes and situations where the proposed change is large, but seldom accepted due to the high energy barrier.
@@ III.5.3 Constant Temperature and Pressure Simulations
III.5.4 Constraints
@@
We implement two constraint algorithms, SETTLE [21] and the constant constraint matrix approximation (CCMA) approach [22]. The former method constraints both positions and velocities by finding analytical expression for three body systems with three constraints, typical for parameterization of water molecules, such as TIP3P [14]. CCMA, on the other hand, is an iterative method that is more generally applicable for arbitrary constraint topologies. It is an iterative Lagrange multiplier-based method, in which the (inverse) of the Jacobian matrix is approximated to be constant by utilizing equilibrium expected behavior. To use constraint algorithms, DIMOS provides a keyword based option for the most common type of constraints when generating the system, i.e., either constraining all bonded interactions containing at least one hydrogen or additionally constraining all angles where at least one of the atoms is a hydrogen.
@@
angle_parameters, tolerance)
III.5.5 Energy minimization
We implement energy minimization by utilizing the optimizer framework available in torch.optim, usually employed for learning network parameters via back-propagation. This allows for easy access to methods like gradient descent or variants thereof including momentum terms, such as commonly used in ML. A maybe more common method for geometry optimization of molecules, also used in OpenMM [26] as standard energy minimizer, is limited-memory BFGS [48]. This is also set as standard in DIMOS.
Having already defined a simulation object, the energy can be minimized very simply by calling
@@ III.5.5 Energy minimization
# If a custom optimizer is chosen:
def optimizer(parameters): return torch.optim.Adam(parameters, lr=lr)
simulation.minimize_energy(steps, optimizer=optimizer)
@@
III.6 Self-tuning Hamiltonian Monte Carlo Simulations
@@
HMC [13] generates MCMC proposals by simulating (discretized) Hamiltonian dynamics. To obtain a new proposal, velocities $v$ are first sampled from the Maxwell-Boltzmann distribution, defining the total energy $E(\mathbf{x},\mathbf{v})=U(\mathbf{x})+T(\mathbf{v})$, where $U(\mathbf{x})$ is the potential energy (dependent only on positions $\mathbf{x}$) and $T(\mathbf{v})$ is the kinetic energy (dependent only on velocities $\mathbf{v}$). In practice, the system is then evolved using a volume-preserving, time-reversible integrator, such as the velocity Verlet integrator. Because phase-space volume is conserved by the integrator, the Metropolis-Hastings acceptance criterion does not require computing the Jacobian determinant.
@@
Short trajectories of only a few integration steps $L$ result in diffusive behavior dominated by random velocity resampling. In contrast, long trajectories in addition to adding additional computational effort risk accumulating numerical errors and redundant exploration (e.g., states “looping back”). Similarly, the integrator timestep $\Delta t$ involves a trade-off: Very small $\Delta t$ lead to inefficient exploration, while large $\Delta t$ introduces effects from a shadow Hamiltonian or even lead to catastrophic instabilities. As shown in prior work [7], naively targeting a fixed acceptance rate (a common heuristic in the field) can fail to optimize the autocorrelation time. This is because acceptance rates only reflect energy conservation errors, not exploration quality.
@@
Instead of tuning the acceptance rate, we optimize HMC simulation parameters via end-to-end training, using a physics-informed objective. In MCMC, a standard optimization metric is the expected squared jump distance (ESJD) [49] of a system property $\theta$, defined as
$\mathcal{L}=\langle(\theta^{\prime}-\theta)^{2}\rangle\approx\frac{1}{R}\sum_{i=0}^{R-1}\alpha_{i}(\theta^{\prime}_{i}-\theta_{i})^{2},$ (15)
@@
$\alpha_{i}=\min(1,\exp-(E(\mathbf{x},\mathbf{v})-E(\mathbf{x^{\prime}},\mathbf{v^{\prime}}))/k_{b}T)$ (16)
depends on the difference of the total energy $E$. Here, $k_{b}$ is the Boltzmann factor, $T$ is the temperature, $\mathbf{x}$,$\mathbf{v}$ are the starting position and (randomly picked) velocity, and $\mathbf{x^{\prime}}$,$\mathbf{v^{\prime}}$ are the proposed state. The average is computed over $R$ independent replicas. For the application in this paper, we choose to optimize directly the movement in real space, i.e., we use $\theta=x$ and set $R=10$.
To optimize $\Delta t$ and $L$, we model them as folded normal variables $\mathcal{N}_{F}(x;\mu,\sigma)$, similar as in Ref. [50], defined on $\mathbb{R}^{+}$ with density:
Here, $\mu$ is learned via backpropagation [50, 51], while $\sigma$ is fixed to $\sigma=0.01\Delta t$ for $\Delta t$ and $\sigma=1$ for $L$. In both cases having $\sigma>0$ functions as a way of jittering, used to avoid effects related to suboptimal performance due to the travelled distance being always the same [7]. For $L$, which is inherently discrete, we reweight contributions to the loss by the step count $L_{j}$, yielding:
$\mathcal{L}=\frac{1}{R}\sum_{i=0}^{R-1}\sum_{j}\frac{w_{j}\alpha_{i,j}}{L_{j}}(\theta^{\prime}_{i,j}-\theta_{i,j})^{2}$ (18)
where $w_{j}=\mathcal{N}_{F}(L_{j};\mu,\sigma)$, and $j$ indexes trajectories up to $L_{j}$. The sum is truncated at $L$, chosen as the quantile function of $\mathcal{N}_{F}$ evaluated at $0.999$. This prioritizes computational efficiency (loss per force evaluation) while avoiding infinite sums. Note that as final proposal we consider the configuration obtained after $\ell$ steps, where $\ell$ is sampled categorically from the discretized distribution of $\mathcal{N}_{F}(\mu,\sigma)$.
We further optimize atomic masses (cf. Fig. 3(d) and (e)): Protein masses are tuned individually, while water oxygen and hydrogen masses are shared across the water molecules. This prevents overfitting to changing solvent configurations. Optimizing the masses is possible in HMC because the partition functions for positions and velocities factorize, hence, the positional distribution is still sampled correctly even though the dynamics does not follow a physical evolution any longer. Since changes in the total mass of the system is trivially equivalent to changes of $\Delta t$, we fix it to be constant and only redistributed masses among the atoms.
To assess the speed-up, we measure the autocorrelation time from the time series of the potential energy defined as
@@
[…]

### v2

@@ II.3 Machine Learning/Molecular Mechanics
An alternative to describing the system fully using an MLIP is to employ a hybrid approach, where part of the system is described using an MLIP, and the remaining part by a classical force field. This is often called ML/MM, standing for machine learning/molecular mechanics, a term derived from QM/MM used to describe the combination of quantum mechanical ab initio calculations and molecular modelling. While the field of QM/MM is quite advanced, in particular in choosing a proper embedding approach for the electrostatic interactions,[37] it is significantly less explored for the ML/MM approach. The current approaches are based on a mechanistic embedding, and as such the electrostatic interactions are treated on the MM level[38] or ignored. To define an ML/MM system, one defines a ml_system which is modeled with an MLIP and a mm_sytem described by a classical force field. In addition, one needs define ml_atoms, a list containing the atom indices modeled by the ml_system, which then results in the new combined system. More details are presented in Section III.2, including example code on how to achieve this is (see Fig. 3(a)).
II.4 Molecular Dynamics and Energy Minimization
@@
With the system description in place, we now present the implemented methods that update the position of the atoms, either to perform molecular dynamics simulations or minimize the potential energy.
II.4.1 Molecular Dynamics
@@
The basis of MD is numerically solving Newtons equation of motion given by $\frac{\partial E}{\partial\vec{r}}=-\vec{F}=-m\vec{a}$. To do so, we implement the ubiquitous methods by Verlet[22] in its velocity variant,[39] which numerically integrates Newtons equation in time and hence conserve the total energy of the shadow Hamiltonian.
@@
Here, $\vec{x}_{i}$ is the position of atom $i$, $\vec{v}_{i}$ and $\vec{a}_{i}$ are the velocity and acceleration. Each step is understood to be performed for all atoms $i$. Since the updates are performed synchronously, the update for different atoms can easily be performed in parallel. In step 3 and 6 the respective constraint algorithms that fix the distances between atoms or angles to their target value is run.
II.4.2 Constant Temperature Simulations
The most direct way to control the temperature of a system, i.e., to sample the canonical ensemble, are thermostats, applied in step 7. For an overview on different thermostatting strategies, see Ref. 40 and references therein. We implement the Berendsen, Andersen, and Lowe-Andersen thermostat.[41] These thermostats work by either rescaling the atoms velocities or by picking new velocities from the Maxwell-Boltzmann distribution for a fraction of atoms or pairs of atoms.[41]
It is also possible to define equations that deviate from Newton’s equation of motion, leading for example to the Langevin and Brownian dynamics equation.[40] These integrators require adapted integration schemes, which we have also implemented in DIMOS.[42]
Another way to control the temperature is to introduce additional degrees of freedom that act as coupling to an external heatbath. This also modifies the integrator, where now the additional degree of freedoms are explicitly integrated using specific splitting schemes. We implement the Nosé–Hoover and Nosé–Hoover chain dynamics,[40] which introduce one or more additional harmonic ‘particles’ that couple the system’s temperature to the target temperature.
Practically, in DIMOS, one defines all different type of thermostats or ways to perform the dynamics as separate integrators. Below, as example, we present the code necessary to define the Langevin and Nosé–Hoover chain dynamics, as well as dynamics using the velocity Verlet integrator in conjunction with the Andersen thermostat.
@@ II.4.2 Constant Temperature Simulations
II.4.3 Constant Temperature and Pressure Simulations
To reproduce many experimental conditions correctly, it is necessary to simulate the system at constant pressure. We implement an isotropic and anisotropic Monte Carlo barostat,[43] which works by proposing new box sizes and accepting the new state with the usual Metropolis Monte Carlo criterion, guaranteeing that the stationary distribution of samples is the target distribution. In particular, one proposes new box sizes for the simulation cell, and then repositions the coordinates of the atoms within this box to correspond to the changed size. A high acceptance rate is achieved by displacing the center of each molecule instead of treating each atom individually, as this way one avoids distortions within each molecule. Alternatively, in cases where molecules are not as clearly defined such as for MLIPs, it is also possible to simply displace every atom independently.
@@ II.4.3 Constant Temperature and Pressure Simulations
Figure 2: Runtime $\tau$ in seconds per MD step for a water box with constant density as a function of number of atoms $N$. (a) Shows the runtimes for torchMD, DIMOS, and openMM using single precision numerics and RFA to model the electrostatic interactions. (b) Demonstrates the influence of different approaches to model the electrostatic interactions: Runtimes are presented for RFA, Ewald, and PME for DIMOS using single precision numerical accuracy. In (c), we investigate the effect of the chosen computational precision on the runtime, modelling electrostatic interactions via RFA in DIMOS. In all cases, the lines indicate the theoretically expected asymptotic computational complexity. (d)-(f) Show the performance of running the MACE foundation model, where in (d) we compare the runtimes of DIMOS to TorchSim, ASE, and LAMMPS using the small MACE model and mixed precision (apart from TorchSim, which uses single precision) for different NN, and (e) investigates the influence of choosing different numerical precision within DIMOS. (f) Presents the runtimes for the biggest system that can be simulated using the large MACE model ($N=8844$) and compares the resulting $\tau$ for DIMOS and ASE.
II.4.4 Bond and Angle Constraints
@@
We implement two constraint algorithms, SETTLE[45] and the constant constraint matrix approximation (CCMA) approach.[46] The former method constraints both positions and velocities by finding analytical expression for three body systems with three constraints, typical for parameterization of water molecules, such as TIP3P.[47] CCMA, on the other hand, is an iterative method that is more generally applicable for arbitrary constraint topologies. It is an iterative Lagrange multiplier-based method, in which the (inverse) of the Jacobian matrix is approximated to be constant by utilizing equilibrium expectation values.
To use constraint algorithms, DIMOS provides a keyword based option for the most common type of constraints when generating the system, i.e., either constraining all bonded interactions containing at least one hydrogen or additionally constraining all angles where at least one of the atoms is a hydrogen.
@@
angle_eq_parameters, tolerance)
II.4.5 Energy minimization
We implement energy minimization by utilizing the optimizer framework available in torch.optim, usually employed for learning network parameters via back-propagation. This allows for easy access to methods like gradient descent or variants thereof including momentum terms, such as commonly used in ML.
By default, we utilize limited-memory BFGS,[48] which is also standard in OpenMM.[3]
Having already defined a simulation object, the energy can be minimized simply by calling
@@ II.4.5 Energy minimization
# If a custom optimizer is used:
def optim(parameters):
return torch.optim.Adam(parameters, lr=lr)
simulation.minimize_energy(steps, optimizer=optim)
@@
III Performance Benchmarks
In the following section, we present the resulting runtimes of DIMOS for simulations of water boxes to probe the influence of system size and simulation/modelling choices such as the method to calculate long-range interactions or different schemes for the numerical datatypes for classical force fields and MLIPs. Then we turn towards the simulation of proteins in explicit solvent, investigating the influence of constraints on the simulation speed and presenting ML/MM as hybrid solution. Finally, we demonstrate how the end-to-end differentiability of the code base allows us to differentiate through the whole simulation to tune parameters of the simulation and system in the framework of HMC.[49, 50]
All simulation results for which we quote runtimes were performed using an NVIDIA RTX 6000 Ada GPU with 48GB memory on a server equipped with an Intel Xeon Gold 5412U CPU with 24 cores and 256GB memory.
ml_atoms = [...] # Indices of atoms that are being modeled by ML system.
[…]
Figure 3: (a) Demonstration on how to set up both classically and machine-learning enhanced systems in DIMOS. ml_atoms| specifies which atoms should be modeled by the MLIP system, and also can be empty in case one does not want to model any systems using ML. (b-d) Achievable simulation time per day $\mathcal{T}$ with and without constrained motion of the hydrogens for each of the three cases of a system modeled purely by ML, the hybrid approach using ML/MM, and a purely MM based modelling. In all cases, the MLIP model was chosen to be mace-mp-large|. In (b), we present the data for alanine-dipeptide (Ala2), (c) shows Trp-Cage, and (d) shows dihydrofolate reductase (DHFR). All systems are modeled using explicit water modeled by the TIP3P parameters.[47] In (d), the biggest system, it is not possible to model purely with ML, as for this system size the MLIP model required more than the available GPU memory.
III.1 System-size Scalability
We investigate the runtimes per simulation step averaged over several simulation steps for water boxes of fixed density with varying number of atoms $N$. In Figs. 2(a)-(c), we present the average runtime per MD step $\tau$ as a function of number of atoms $N$ in an unconstrained simulation of water under Langevin dynamics at $300$ K with timestep $\Delta t=0.5$ fs, friction coefficient $\gamma=0.1$ fs-1, and cutoff $r_{c}=10$ Å. The water is modeled by the TIP3P water model.[47]
In (a) we compare $\tau$ obtained in DIMOS to torchMD[9] and openMM[3] for simulations running in single floating point precision and electrostatic interactions modeled by the reaction field approximation (the RFA is the only setup for electrostatics supported by torchMD). We observe linear scaling of the runtime per step $t$ in DIMOS and openMM, whereas torchMD displays quadratic scaling as a function of number of atoms $N$. For the largest system size tested ($\approx 300,000$ atoms), DIMOS as a result achieves a $170\times$ speed-up over torchMD, but is factor $40\times$ slower than openMM. For torchMD, the runtime of the largest system size is extrapolated from the expected quadratic complexity, while, as expected, DIMOS and openMM have linear scaling due to utilizing neighborlists and cell lists. [^3] For a fair comparison, we check the validity of the neighborlist after every integration step, avoiding any potential artifacts leading to inaccuracies in the force/energy evaluation. The slowdown compared to openMM is epected, since DIMOS is implemented on a much higher abstraction level prohibiting many optimizations. This aligns with our envisioned role of DIMOS as a complementary platform for rapid prototyping and gradient-based workflows rather than a direct replacement for production-optimized codes
Panel (b) demonstrates the influence of different methods how to model electrostatics within DIMOS. We use single precision numerics. Clearly the RFA is the fastest option due to the simple mean-field treatment of charges outside the cutoff. Somewhat slower, but still competitive is PME,[29] for which one expects $O(N\log N)$ complexity, which roughly agrees with our data. In contrast, using direct Ewald summation[28] with theoretically ideal complexity $O(N^{3/2})$ is significantly slower already for medium-sized systems, and can only provide a minor performance improvement over PME for small systems. In addition, direct Ewald summation requires significantly more GPU memory and can thus not be used for large systems.
The influence of the chosen numerical accuracy is presented in Fig. 2(c) with electrostatic interactions modeled by RFA, where we find that double precision as expected is overall slower on our GPU. Very little performance penalty is observed using mixed precision, where the integration is performed in double precision, but the force/energy evaluation is performed in single precision. However, this approach cannot be used for larger systems, as the memory requirements increase significantly. If memory consumption is not a bottleneck, it is advisable to always consider using mixed precision due to the reduction of accumulation of errors in comparison to single precision, where such effects can be detrimental.
In Figs. 2(d)-(f) we investigate the behavior of DIMOS in conjunction with MLIPs by simulating the same water boxes as in (a)-(c) with the mace-mp foundational model.[6, 35] MACE is a state-of-the-art equivariant message-passing graph neural network that models interatomic interactions. The simulation parameters remain largely the same, however, we now employ the standard interaction cutoff for message passing of $r_{c}=3$ Å in MACE.
In panel (d), we compare DIMOS to three different simulation engines: i) TorchSim,[12] a (differentiable) torch-based simulation package released after DIMOS focusing on batching simulations, ii) ASE,[52] one of the most popular (non-differentiable) approaches to run MLIP based simulation due to its simplicity, and iii) LAMMPS,[2] using the ML-IAP interface as example of a compiled simulation engine, prioritizing simulation speed but, as discussed previously, incurring additional overhead in model setup. Apart from TorchSim, for which we use single precision since mixed precision is not supported, all simulations are performed in mixed precision. We observe that for the smallest foundation model mace-mp-small, DIMOS positions itself as the fastest high-level simulation engine. TorchSim suffers from slow simulation speed and large memory demands, the latter limiting the maximal system size we can simulate. [^4] We follow the official TorchSim tutorial to set up the simulation. As a consequence, for the biggest system size that is simulated by TorchSim of $N=6282$, DIMOS is roughly factor $\approx 5$ faster than TorchSim. DIMOS is also faster than ASE by factor $\approx 1.7$ for the biggest system size $N=12255$. Compared to LAMMPS, DIMOS is roughly $2.5\times$ slower, but employing LAMMPS, as discussed, comes with constraints on the atomic descriptor and requires the model to be implemented using TorchScript. Consequently, for large-scale simulations, it may be advantageous to adopt a hybrid approach that leverages the respective strengths of both packages to maximize performance and flexibility.
In (e), we present the influence of choosing different numerical precision, i.e., using single, mixed, or double precision. There is virtually no penalty in using mixed precision, because the runtime and memory consumption is dominated by the ML model, so that all other factors effectively become negligible, at least for the currently achievable system sizes.
Figure 2(f) shows the influence of choosing different sizes of the mace-mp model in comparison between DIMOS and ASE for the biggest system size ($N=8844$) that can be simulated using all ML model sizes. We here focus on ASE, since TorchSim has higher memory requirements and slower simulation speed and LAMMPS is less directly comparable in user experience. For the small model, the speed-up obtained by DIMOS is the biggest with factor $1.6\times$, i.e., nearly half of the time in ASE is spent in integrating the equation of motion and memory transfer. For the bigger models, the difference between DIMOS and ASE become less pronounced due to the larger time spent in the actual ML model evaluation, but the speed-up still remains sizable.
III.2 Constraints and Hybrid ML/MM for Protein Systems
We simulate three protein benchmark systems with explicit solvent of varying size: i) Alaninedipeptide (Ala2) with $2269$ atoms, ii) Trp-Cage with $7660$ atoms, and iii) dihydrofolate reductase (DHFR) with $23558$ atoms. The first two systems are standard test-systems to probe novel simulation methods, and the latter is a common testsystem from the joint AMBER/CHARMM benchmark. [^5] See https://ambermd.org/GPUPerformance.php. Since we here also want to probe the effect of using constraint algorithms to increase the timestep on the achievable simulation time, we shift our metric to the simulation time achievable per wall-clock day $\mathcal{T}$, i.e., how many ns of simulation can be achieved per day. In the simulations we use the Nosé–Hoover integrator with collision frequency of $0.05$ fs-1 and $\Delta t=0.5$ fs timestep for disabled constraints, and $\Delta t=2.0$ fs timestep with constraints enabled.
We are considering three scenarios: i) The system purely modeled by mace-mp-large (ML), ii) a hybrid approach where the protein is modeled by the MACE model (ML/MM), but the water is parameterized by the classical TIP3P water model,[47] and iii) a fully classical simulation using AMBER99SB[18] force field and TIP3P water (MM).[47] In the ML and ML/MM case, electrostatic contributions are ignored (DIMOS implements several options), whereas in the MM case we use PME. As demonstrated in Fig. 3(a), it is straightforward to set up systems modeled by the three options in DIMOS. The only option to be specified by the user is which MLIP to use, and which atoms should be modeled by it.
In Fig. 3(b), we present the data for Ala2 using the three approaches with and without constraints. For the systems modeled by ML or ML/MM, we adopt the constraints as defined for the MM system. The ML approach is clearly the slowest option with $\mathcal{T}\approx 0.18$ ns/day of simulation time without constraints. This is in contrast to the fastest simulation time $\mathcal{T}\approx 20$ ns/day (more than $100\times$ faster than the ML approach), achieved by modelling the interaction using MM and with constraints activated. As viable alternative, one may consider the hybrid approach, where with constraints we achieve $\mathcal{T}\approx 5.4$ ns/day.
For the larger Trp-Cage system presented in (c), the differences become even more pronounced. While a classical ML simulation without constraints achieves $\mathcal{T}\approx 0.05$ ns/day, the ML/MM approach without constraints is $14\times$ faster. In the extreme case, employing constraints and modelling all interactions using MM, we observe a speed-up of $175\times$. While modelling a part or all of the interactions using MM comes with an accuracy penalty, that this is easily possible to be implemented in DIMOS highlights the modularity of the framework with easy combination of various techniques tailored to the problem.
Figure 3(d) shows the performance for DHFR, which due to memory constraints of MACE could not be simulated completely using the ML approach on our hardware. Using ML/MM, we obtain runtimes in the range of $\approx 0.15$ to $\mathcal{T}\approx 0.6$ ns/day, and $\approx 3.8$ to $\approx 6.5$ ns/day for pure MM modelling.
IV End-to-end Optimization of Hamiltonian Monte Carlo
Figure 4: Plots demonstrating the end-to-end learning of simulation parameters for HMC. (a) Loss $\mathcal{L}_{\mathrm{HMC}}$ as function of training time $t$. (b),(c) The mean of timestep $\mu_{\Delta t}$ and number of integration steps $\mu_{L}$ against training time $t$. (d),(e) The behavior of the masses $m$ for (d) the hydrogen and oxygen atoms in the water molecules, optimized as shared parameter between all water molecules and (e) the remaining masses in Ala2. In (e) we plot $m-m_{0}$ to better highlight the relative change of masses for the diverse atom types. (f) Autocorrelation function $A(k)$ of the potential energy $U(x)$ as function of lag-time $k$ measured in terms of the number of energy evaluations.
In DIMOS, a MCMC simulation is set up by defining the chosen proposals, which are then together with the system used to initialize the simulation object:
hmc_move = dimos.monte_carlo.HMC(T, dt, L, system)
In above example, we define a standard HMC move and assign it probability $1.0$ of being picked as proposal distribution when initializing the simulation object.
It is possible to associate several types of MCMC move sets to generate a simulation object, where the different moves are picked with the probabilities specified in the associated list. Here, we are interested in HMC moves, and to tune its parameters such that the proposal distribution is optimized to reduce the autocorrelation between samples. For this, we utilize the end-to-end differentiability of DIMOS. Other packages that in principle could be used are TorchMD[9] and TorchSim,[12] or other Jax or Julia-based simulation engines such as Jax M.D.,[8] FeNNol,[11] or Molly.jl.[10] In our original work where we optimized HMC for toy systems and a small protein in vacuum by end-to-end differentation,[17] we indeed relied on TorchMD.[9] However, as presented above, for larger systems this would incur a significant computational overhead.
@@ IV End-to-end Optimization of Hamiltonian Monte Carlo
HMC[49, 50] constructs proposal configurations that are informed by the forces acting on the particles, in contrast to random-walk MCMC where proposals are uninformed. This allows for large configurational changes without incurring a low acceptance rate. To obtain a new proposal in HMC, the configuration space is extended to include velocities $v$ (which is usually also done in molecular dynamics simulations). A proposal is then generated by first sampling velocities $v$ from the Maxwell-Boltzmann distribution, defining the total energy $E(\mathbf{x},\mathbf{v})=U(\mathbf{x})+T(\mathbf{v})$, where $U(\mathbf{x})$ is the potential energy (dependent only on positions $\mathbf{x}$) and $T(\mathbf{v})$ is the kinetic energy (dependent only on velocities $\mathbf{v}$). In practice, the system is then evolved using a volume-preserving, time-reversible integrator, such as the velocity Verlet integrator of Section II.4.1. Because phase-space volume of the extended system is conserved by the integrator, the Metropolis-Hastings acceptance criterion does not require computing the Jacobian determinant.
@@
Short trajectories of only a few integration steps $L$ result in diffusive behavior dominated by random velocity resampling. In contrast, long trajectories in addition to adding additional computational effort risk accumulating numerical errors and redundant exploration (i.e., states “looping back”). Similarly, the integrator timestep $\Delta t$ involves a trade-off: Small $\Delta t$ lead to inefficient exploration, while large $\Delta t$ introduces effects from a shadow Hamiltonian or even lead to catastrophic instabilities. As shown in prior work,[17] naively targeting a fixed acceptance rate can fail to optimize the autocorrelation time under certain conditions. This is because acceptance rates only reflect energy conservation errors, not exploration quality.
@@
Instead of tuning the acceptance rate, we optimize HMC simulation parameters via end-to-end training, using a physics-informed objective. Our parameters are the timestep $\Delta t$, the number of integration steps $L$, and the masses of the atoms $m_{i}$. We optimize the expected squared jump distance (ESJD)[55] of a system observable $\mathcal{O}$, defined as
$\mathcal{L}=\langle(\mathcal{O}^{\prime}-\mathcal{O})^{2}\rangle\approx\frac{1}{R}\sum_{i=1}^{R}\alpha_{i}(\mathcal{O}^{\prime}_{i}-\mathcal{O}_{i})^{2},$ (15)
@@
$\alpha_{i}=\min(1,\exp(-[E(\mathbf{x},\mathbf{v})-E(\mathbf{x^{\prime}},\mathbf{v^{\prime}})]/k_{b}T))$ (16)
depends on the difference of the total energy $E$. Here, $k_{b}$ is the Boltzmann factor, $T$ is the temperature, $\mathbf{x}$$\mathbf{v}$ are the starting position and (randomly picked) velocity, and $\mathbf{x^{\prime}}$$\mathbf{v^{\prime}}$ is the proposed state. The average is computed over $R$ independent replicas. For the application in this paper, we set $R=10$ and choose to optimize directly the movement in real space, i.e., we use $\mathcal{O}=x$.
This loss is used as a proxy for the autocorrelation time, defined as
@@
where $\langle\ldots\rangle$ symbolizes the thermodynamic expectation and $k$ is the lag-time. For large $k$, $A(k)$ decays exponentially